OpenStudy (anonymous):

Please help me with this identity: (cosx+sinx)(1-sinxcosx)=cos ^{3}x+sin ^{3}x

OpenStudy (anonymous):

(cosx+sinx)(1-sinxcosx)=cos ^{3}x+sin ^{3}x

OpenStudy (anonymous):

\[(cosx+sinx)(1-sinxcosx)=\cos ^{3}x+\sin ^{3}x\]

OpenStudy (zehanz):

You will need this factorisation: a³+b³ = (a+b)(a²-ab+b²). You can verify it by doing a long division of a³+b³ by a+b. Because this is a trig problem, I don't think you are required to do that here, so: \(\cos^3x+\sin^3x=\(\cos x+\sin x)(\cos^2x - \sin x\cos x+ \sin^2x)\). If you compare this reworked right-hand side with the LHS, I think you are almost done...

OpenStudy (zehanz):

(ignore the red "\" above...)

OpenStudy (zehanz):

Maybe it could also be done without first looking at the RHS. If you replace 1 with sin²x+cos²x in the LHS, if you work out all the multiplications (6 in total) you should eventually arrive at the RHS...

OpenStudy (anonymous):

thanks, yeah i tried multiplying out the left side but didn't replace the one and got stuck. Using the sum of 2 cubes trick got me to the right answer. Thank you for all your help today

OpenStudy (zehanz):

No problem!